• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name yez

    Log entry time 18:15:12 on February 02, 2011

    Entry number 346049

    keyword=L-HRS: Cosmic Ray run restuls update

    I analyzed the cosmic run we took in this week. I found several problems
    and tried to fix them. Here is the details:
    1, VDC,PR1,PR2 are all working fine
    2, S1.ADC.#1R is empty because it was connected to a scope but not ADC
    module.  *FIXED*
    3, S1.TDC.#1R/#2R/#3R are empty because the connector of the cable to TDC
    module is fliped. *FIXED*
    4, S2m.ADC.#10L~#16L, the connector was loose. *FIXED*
    5, Cer-TDC, all channels have no peaks, because the signals were not
    delayed before they being sent into 1875TDC. We will put the cable into
    6, S2m.TDC.#3R is empty since the Level Translator (Crate#2,Slot#5) has a
    bad channels at #3. Need to replace the Level Translator or fix it.

    A copy of this log entry has been emailed to: brads@jlab.org,vasulk@jlab.org

    Figure 1

    Figure 2

    Figure 3

    Figure 4

    Figure 5

    Figure 6

    Figure 7

    Figure 8

    Figure 9

    Figure 10